# Writing – original draft, Data curation, Formal analysis, Methodology, Conceptualization

### v1

@@ Abstract
Preterm birth (PTB), defined as delivery before 37 weeks of gestation, is a leading cause of neonatal mortality and long term health complications. Early detection is essential for enabling timely medical interventions. Electrohysterography (EHG) and tocography (TOCO) are promising non invasive tools for PTB prediction, but prior studies often suffer from class imbalance, improper oversampling, and reliance on features with limited physiological relevance. This work presents a machine learning pipeline incorporating robust preprocessing, physiologically grounded feature extraction, and rigorous evaluation. Features were extracted from EHG (and TOCO) signals using Mel frequency cepstral coefficients, statistical descriptors of wavelet coefficients, and peaks of the normalized power spectrum. Signal quality was enhanced via Karhunen Loève Transform (KLT) denoising through eigenvalue based subspace decomposition. Multiple classifiers, including Logistic Regression, Support Vector Machines, Random Forest, Gradient Boosting, Multilayer Perceptron, and CatBoost, were evaluated on the TPEHGT dataset. The CatBoost classifier with KLT denoising achieved the highest performance on fixed interval segments of the TPEHGT dataset, reaching 97.28% accuracy and an AUC of 0.9988. Ablation studies confirmed the critical role of both KLT denoising and physiologically informed features. Comparative analysis showed that including TOCO signals did not substantially improve prediction over EHG alone, highlighting the sufficiency of EHG for PTB detection. These results demonstrate that combining denoising with domain relevant features can yield highly accurate, robust, and clinically interpretable models, supporting the development of cost effective and accessible PTB prediction tools, particularly in low resource healthcare settings.
keywords:
@@ 1 Introduction
To overcome data imbalance, many researchers apply different augmentation techniques [7]. However, as Vandewiele et al. [14] pointed out, most high performing benchmark models applied oversampling before splitting the dataset into training and testing sets. This approach inflates performance metrics, making the results unreliable. Vandewiele et al. highlighted this by revisiting the original study [12] behind the widely used public TPEHGT dataset. The study reported an AUC of 99.44 using quadratic discriminant analysis (QDA) with adaptive synthetic (ADASYN) oversampling, but when oversampling was correctly applied after data splitting, performance dropped to 62.36. Their findings exposed major flaws in past studies that adopted these evaluation practices, constituting the majority of the work, emphasizing the need for more rigorous and realistic modeling in this critical domain.
@@
From a physiological perspective, specific frequency bands in EHG signals are known to reflect meaningful uterine and maternal activity. The 0.08–1.0 Hz range is primarily linked to uterine physiological activity, while the 1.0–2.2 Hz, 2.2–3.5 Hz, and 3.5–5.0 Hz bands are associated with the maternal heart’s influence on the uterus, including the heart rate and its second and third harmonics [16]. Notably, the 1.0–2.2 Hz band, which reflects maternal heart induced electrical and mechanical uterine activity, tends to show higher amplitude during term pregnancies when labor is still distant, and lower amplitude as delivery approaches. Interestingly, this band is also observed to have low amplitudes during preterm pregnancies, even when delivery is not expected soon [12]. This behavior suggests that specific sub bands in the frequency domain carry predictive information, emphasizing the need for targeted feature extraction to capture the underlying physiological signals linked to preterm birth.
In this paper, we revisit existing publicly available datasets and address key gaps related to improper oversampling and the limited focus on extracting physiologically meaningful frequency domain features from EHG signals. Our approach introduces a reliable and reproducible pipeline for PTB detection, validated on a second dataset to demonstrate robustness. The main contributions of this work are:
@@
To present our study clearly, the remainder of this paper is organized as follows. Section 2 describes the datasets, preprocessing steps, feature extraction, and classification models used in the proposed approach. Section 3 presents the experimental results and discussion, compares them with existing methods, and includes an ablation study. Finally, Section 4 summarizes the main contributions of this work.
2 Material and methods
@@
2.1 Dataset
This study employs two publicly available EHG datasets hosted on PhysioNet: the TPEHGT dataset, published in 2018 [12], and the TPEHG dataset, published in 2012 [13]. Both datasets were developed at the Faculty of Computer and Information Science, University of Ljubljana, Slovenia, with recordings collected at the Department of Obstetrics and Gynecology, University Medical Center Ljubljana. The main experiments and results presented in the following sections were conducted using the widely adopted TPEHGT dataset. To assess the robustness of our work, we additionally evaluated it on the TPEHG dataset.
@@ 2.1 Dataset
2.2 Preprocessing
2.2.1 Band-pass filtering
@@
To do this, we applied a coarse Butterworth bandpass filter in the 0.08–5 Hz range to remove outright noise. Fig. 2 illustrates the PSD of the signal before and after filtering. The filtered signal was then used for spectral feature engineering. Subsequently, finer signal subspace separation and feature selection were performed using KLT as mentioned in Section 2.2.2. This hierarchical approach avoids biasing the system toward predefined bands, allowing the model to learn the most predictive frequency domain characteristics and justifying the use of KLT as noted in our contributions.
@@
2.2.2 Karhunen Loève Transform
While band pass filtering reduces broad spectrum noise and isolates physiologically relevant frequency components, it does not fully eliminate all sources of noise or artifacts in the EHG signal. In particular, residual noise and overlapping components may remain within the filtered bandwidth, potentially degrading the quality of the features. To further refine the signal and enhance its structural clarity, we applied the KLT to remove noise components.
The KLT decomposes the signal into an orthogonal basis derived from the data’s covariance structure, where each component (eigenvector) captures a portion of the signal’s variance, which in this context corresponds to energy. The associated eigenvalues quantify the energy captured along each eigenvector direction, with the largest eigenvalues representing the dominant signal components. When the covariance matrix $R$ is approximately Toeplitz, these eigenvector directions align with the spectral context of the signal under the assumption of reasonable stationarity. In contrast, white noise, being random and uncorrelated, distributes its energy uniformly across directions, leading to smaller and more evenly spread eigenvalues. These smaller eigenvalues form the noise subspace, while the dominant signal remains concentrated in the eigenvectors associated with the largest eigenvalues. This distinction underpins our selection criteria for separating signal from noise.
By retaining only the components associated with the largest eigenvalues and discarding those with small and evenly spread eigenvalues (corresponding to noise), we effectively separate and suppress the noise dominated subspace while preserving the meaningful structure of the signal.
Let $x[n]$ denote an input discrete time signal of length $N$. To compute the KLT, we first estimate the autocorrelation matrix $R$ assuming a wide sense stationary signal in the selected small window. Rather than computing the full $N\times N$ matrix, we restrict the autocorrelation to a fixed lag, $L=50$ to focus on local temporal structure and reduce computational complexity. The autocorrelation sequence $r[k]$ is estimated for a fixed lag $L$ using:
$r[k]=\frac{1}{N-k}\sum_{n=0}^{N-k-1}x[n]\cdot x[n+k],\quad\text{for }k=0,1,\dots,L-1.$ (1)
This provides an unbiased estimate of the autocorrelation sequence up to lag L1L-1. The autocorrelation matrix $R\in\mathbb{R}^{L\times L}$ is then constructed as a symmetric Toeplitz matrix:
$R=\begin{bmatrix}r[0]&r[1]&r[2]&\cdots&r[L-1]\\ r[1]&r[0]&r[1]&\cdots&r[L-2]\\ r[2]&r[1]&r[0]&\cdots&r[L-3]\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ r[L-1]&r[L-2]&r[L-3]&\cdots&r[0]\end{bmatrix}$ (2)
The KLT basis is derived by decomposing $R$ into its eigenvalues and eigenvectors:
@@ 2.2.2 Karhunen Loève Transform
where $\Lambda$ is a diagonal matrix of eigenvalues sorted from smallest to largest ($\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{N}$), and $Q$ contains the corresponding eigenvectors.
To identify the significant eigenvectors, we first compute the logarithm of the sorted eigenvalues ($\log\lambda_{i}$) and then calculate the relative difference between consecutive log-eigenvalues. A sharp increase in this difference indicates a transition from signal dominated components to noise dominated ones. In our analysis, we define this transition point as the first instance where the relative change exceeds 10%, as illustrated in Fig. 3. All eigenvectors $Q_{k},Q_{k+1},\dots,Q_{N}$ associated with eigenvalues above this threshold ($\lambda_{k+i}>\lambda_{\text{threshold}}$) are retained for reconstructing the denoised signal.
@@
Figure 3: (a) Log-eigenvalue spectrum ($\log\lambda_{i}$) in ascending order. (b) Relative inter eigenvalue differences obtained by subtracting consecutive points. The threshold marks a 10% increase, separating noise dominated eigenvalues (left) from signal dominated components (right) retained for reconstruction.
The signal is then projected onto the eigenvectors corresponding to the dominant signal subspace:
$X^{\prime}=Q_{k}^{T}X$ (4)
where $X$ is the zero mean input signal and $Q_{k}$ spans the selected eigenvectors subspace. This projection effectively decorrelates the signal into multiple sub components aligned with the directions of the signal subspace.
The denoised signal $\widehat{X}$ is then reconstructed using the inverse transform:
$\widehat{X}=Q_{k}X^{\prime}.$ (5)
@@
This reconstruction suppresses contributions from the noise subspace (smaller, uniformly distributed eigenvalues) and retains only the dominant components of the signal. The resulting noise filtered signal is then used for feature extraction and subsequent classification.
2.3 Feature Extraction
2.3.1 Peak amplitude of the normalized power spectrum
The PA of the normalized power spectrum quantifies the dominant frequency component in a specified band, normalized by the total power in that band. This feature helps distinguish between term and preterm uterine activity.
@@ 2.3.1 Peak amplitude of the normalized power spectrum
For a discrete signal $x[n]$, the power spectral density (PSD) $P(f_{i})$ is estimated over discrete frequency bins $f_{i}$ within a frequency band $f_{\text{low}},f_{\text{high}}]$ using Welch’s method. The $P(f_{i})$ of the EHG signal was then normalized so that the total spectral power equals 1, which gives compatibility across all the different EHG recordings. In the given band of normalized EHG signal, the highest value was selected as the PA of the normalized power spectrum.
$A_{\text{peak}}=\max_{f_{i}\in[f_{\text{low}},f_{\text{high}}]}\frac{P(f_{i})}{\sum_{f_{j}=f_{\text{low}}}^{f_{\text{high}}}P(f_{j})}$ (6)
2.3.2 Mel coefficients extraction
Mel Frequency Cepstral Coefficients (MFCCs) represent the short term power spectrum of a signal and are commonly used for feature extraction based on the Short-Time Fourier Transform (STFT). Initially developed for speech recognition, MFCCs have gained attention in biomedical signal processing due to their ability to capture perceptually relevant spectral features. MFCCs employ a Mel scale filter bank, which maps the frequency axis nonlinearly, providing higher resolution for lower frequencies. This emphasis is beneficial since both biomedical and acoustic signals are highly dependent on information present in the low frequency range [17].
@@ 2.3.2 Mel coefficients extraction
2.3.3 Wavelet coefficients extraction
@@
In this work, the Daubechies 8 (db8) wavelet was selected because it effectively captures both the smooth low frequency uterine activity and the sharp transient changes, making it more suitable than shorter Daubechies wavelets. This property makes it particularly appropriate for analyzing the nonstationary nature of EHG signals [21]. Accordingly, the DWT was implemented using the db8 wavelet with a 5-level decomposition to analyze the EHG signals.
For each channel of the EHG signals, the 5-level wavelet decomposition produces five sets of detail coefficients (D1-D5) and one set of approximation coefficients (A5), yielding six sub bands in total. From each sub band, six statistical features were extracted from the wavelet coefficients to characterize the signal’s time-frequency properties. This results in 6 features $\times$ 6 sub-bands = 36 features per channel. These features are computed from the coefficients within each sub band, where $c[i]$ represents the $i$-th wavelet coefficient and $N$ is the total number of coefficients in that sub band.
@@
2.4 Model Selection
Given the characteristics of the dataset and the relatively limited size of available labeled EHG recordings, classical ML techniques were employed for classifying signal segments. These standard ML methods were chosen because, with limited data, more complex approaches such as DL can easily overfit. Coupled with the efficient denoising and feature selection strategies described in the previous sections, these computationally simpler methods are sufficient to capture relevant patterns. A set of supervised binary classification models was selected based on their proven performance in prior studies focused on PTB prediction and similar biomedical signal classification tasks [8, 7]. The selected models include:

### v2

@@ Abstract
Preterm birth (PTB), defined as delivery before 37 weeks of gestation, is a leading cause of neonatal mortality and long term health complications. Early detection is essential for enabling timely medical interventions. Electrohysterography (EHG) and tocography (TOCO) are promising non invasive tools for PTB prediction, but prior studies often suffer from class imbalance, improper oversampling, and reliance on features with limited physiological relevance. This work presents a machine learning (ML) pipeline incorporating robust preprocessing, physiologically grounded feature extraction, and rigorous evaluation. Features were extracted from EHG (and TOCO) signals using Mel frequency cepstral coefficients, statistical descriptors of wavelet coefficients, and peaks of the normalized power spectrum. Signal quality was enhanced via Karhunen Loève Transform (KLT) denoising through eigenvalue based subspace decomposition. Multiple classifiers, including Logistic Regression, Support Vector Machines, Random Forest, Gradient Boosting, Multilayer Perceptron, and CatBoost, were evaluated on the TPEHGT dataset. The CatBoost classifier with KLT denoising achieved the highest performance on fixed interval segments of the TPEHGT dataset, reaching 97.28% accuracy and an AUC of 0.9988. Ablation studies confirmed the critical role of both KLT denoising and physiologically informed features. Comparative analysis showed that including TOCO signals did not substantially improve prediction over EHG alone, highlighting the sufficiency of EHG for PTB detection. These results demonstrate that combining denoising with domain-relevant features can yield highly accurate, robust, and clinically interpretable models, supporting the development of cost-effective and accessible PTB prediction tools, particularly in low-resource healthcare settings.
Index Terms:
Electrohysterography (EHG), Preterm birth detection, Uterine contractions, Biomedical signal analysis, Machine learning (ML)
@@ I Introduction
To overcome data imbalance, many researchers apply different augmentation techniques [7]. However, as Vandewiele et al. [14] pointed out, most high performing benchmark models applied oversampling [15] before splitting the dataset into training and testing sets. This approach inflates performance metrics, making the results unreliable. Vandewiele et al. highlighted this by revisiting the original study [12] behind the widely used public TPEHGT dataset. The study reported an AUC of 99.44 using quadratic discriminant analysis (QDA) with adaptive synthetic (ADASYN) oversampling, but when oversampling was correctly applied after data splitting, performance dropped to 62.36. Their findings exposed major flaws in past studies that adopted these evaluation practices, constituting the majority of the work, emphasizing the need for more rigorous and realistic modeling in this critical domain.
@@
From a physiological perspective, specific frequency bands in EHG signals are known to reflect meaningful uterine and maternal activity. The 0.08–1.0 Hz range is primarily linked to uterine physiological activity, while the 1.0–2.2 Hz, 2.2–3.5 Hz, and 3.5–5.0 Hz bands are associated with the maternal heart’s influence on the uterus, including the heart rate and its second and third harmonics [17]. Notably, the 1.0–2.2 Hz band, which reflects maternal heart induced electrical and mechanical uterine activity, tends to show higher amplitude during term pregnancies when labor is still distant, and lower amplitude as delivery approaches. Interestingly, this band is also observed to have low amplitudes during preterm pregnancies, even when delivery is not expected soon [12]. This behavior suggests that specific sub bands in the frequency domain carry predictive information, emphasizing the need for targeted feature extraction to capture the underlying physiological signals linked to PTB.
In this paper, we revisit existing publicly available datasets and address key gaps related to improper oversampling and the limited focus on extracting physiologically meaningful frequency domain features from EHG signals. Our approach introduces a reliable and reproducible pipeline for PTB detection, evaluated on a second publicly available dataset to assess consistency. The main contributions of this work are:
@@
To present our study clearly, the remainder of this paper is organized as follows. Section II describes the datasets, preprocessing steps, feature extraction, and classification models used in the proposed approach. Section III presents the experimental results and discussion, compares them with existing methods, and includes an ablation study. Finally, Section IV summarizes the main contributions of this work.
II Materials and Methods
@@
II-A Dataset
This study employs two publicly available EHG datasets hosted on PhysioNet: the TPEHGT dataset, published in 2018 [12], and the TPEHG dataset, published in 2012 [13]. Both datasets were developed at the Faculty of Computer and Information Science, University of Ljubljana, Slovenia, with recordings collected at the Department of Obstetrics and Gynecology, University Medical Center Ljubljana. The main experiments and results presented in the following sections were conducted using the widely adopted TPEHGT dataset. To further evaluate the proposed pipeline on an additional dataset, we report results on the TPEHG dataset.
@@ II-A Dataset
II-B Preprocessing
II-B1 Band-pass filtering
@@
To do this, we applied a coarse Butterworth bandpass filter in the 0.08–5 Hz range to remove outright noise. Fig. 2 illustrates the PSD of the signal before and after filtering. The filtered signal was then used for spectral feature engineering. Subsequently, finer signal subspace separation and feature selection were performed using KLT as mentioned in Section II-B2. This hierarchical approach avoids biasing the system toward predefined bands, allowing the model to learn the most predictive frequency domain characteristics and justifying the use of KLT as noted in our contributions.
@@
II-B2 Karhunen Loève Transform
While band pass filtering suppresses out of band interference and isolates physiologically relevant frequency components, residual in band noise and artifacts may still remain and can degrade feature quality. To further refine the EHG signal, we apply a KLT based subspace projection to attenuate noise dominated components while preserving the structured content of the signal.
The KLT represents a signal in an orthogonal basis obtained from the eigendecomposition of its second order statistics. For a wide sense stationary process, the autocorrelation matrix is approximately Toeplitz, and its eigenvectors capture dominant correlation patterns. In contrast, white noise is approximately uncorrelated, leading to a covariance close to $\sigma^{2}I$ with a relatively flat eigenvalue spectrum (i.e., no dominant directions). Motivated by this behavior, we separate a signal dominated subspace associated with larger eigenvalues from a noise dominated subspace associated with small, weakly varying eigenvalues.
Let $x[n]$ denote an input discrete-time signal of length $N$. Rather than forming the full $N\times N$ correlation matrix, we restrict the autocorrelation to a fixed lag $L=50$ to focus on local temporal structure and reduce computational complexity. The unbiased autocorrelation sequence is estimated as
$r[k]=\frac{1}{N-k}\sum_{n=0}^{N-k-1}x[n]\,x[n+k],\quad k=0,1,\dots,L-1.$ (1)
Using the first $L$ lags, the autocorrelation matrix $R\in\mathbb{R}^{L\times L}$ is constructed as a symmetric Toeplitz matrix:
$R=\begin{bmatrix}r[0]&r[1]&r[2]&\cdots&r[L-1]\\ r[1]&r[0]&r[1]&\cdots&r[L-2]\\ r[2]&r[1]&r[0]&\cdots&r[L-3]\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ r[L-1]&r[L-2]&r[L-3]&\cdots&r[0]\end{bmatrix}$. (2)
The KLT basis is obtained from the eigendecomposition of $R$:
@@ II-B2 Karhunen Loève Transform
where $\Lambda=\mathrm{diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{L})$ contains eigenvalues and $Q=[q_{1},q_{2},\dots,q_{L}]$ contains the corresponding eigenvectors. Eigenvalues are sorted in ascending order ($\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{L}$). In implementation, eigenvectors are $\ell_{2}$ normalized to unit norm to form a well defined orthogonal projection.
Subspace selection via a 10% eigen jump: To identify the transition between noise dominated and signal dominated components, we analyze the differences between consecutive log eigenvalues:
$\Delta_{i}=\log(\tilde{\lambda}_{i+1})-\log(\tilde{\lambda}_{i}),\quad i=1,2,\dots,L-1,$ (4)
where $\tilde{\lambda}_{i}=|\lambda_{i}|+\epsilon$ (with a small $\epsilon$) is used for numerical stability. We define the transition index $k$ as the first instance where the log eigenvalue jump exceeds a 10% criterion,
$\Delta_{k}>\tau,\quad\tau=0.1,$ (5)
which corresponds to approximately a 10% increase in eigenvalue magnitude (since $e^{0.1}\approx 1.105$). If no such index exists, $k$ is set to the index of the maximum jump. Fig. 3 illustrates the log eigenvalue spectrum and the corresponding inter-eigenvalue differences. The eigenvectors associated with the larger eigenvalues beyond this transition are retained as the signal subspace:
$Q_{s}=[q_{k+1},q_{k+2},\dots,q_{L}].$ (6)
@@
Figure 3: (a) Log-eigenvalue spectrum $\log(\lambda_{i})$ in ascending order. (b) Consecutive log-eigenvalue differences $\Delta_{i}$. The 10% jump threshold ($\tau=0.1$) identifies the transition from noise-dominated components (left) to signal-dominated components (right) retained for reconstruction.
Block wise denoising via subspace projection: The denoised signal is obtained by projecting the input onto the selected subspace. Specifically, the signal is partitioned into non overlapping blocks of length $L$, forming vectors $\mathbf{x}_{m}\in\mathbb{R}^{L}$. Each block is denoised using the orthogonal projector $Q_{s}Q_{s}^{T}$:
$\widehat{\mathbf{x}}_{m}=Q_{s}Q_{s}^{T}\mathbf{x}_{m}.$ (7)
@@
The final denoised signal $\widehat{x}[n]$ is obtained by concatenating the reconstructed blocks. This procedure suppresses contributions from the noise dominated subspace while retaining the dominant correlated structure of the EHG signal. The resulting noise filtered signal is then used for feature extraction and subsequent classification.
II-C Segmentation and Sample Construction
This work evaluates two segmentation strategies that define the sample unit used for feature extraction and classification: (i) contraction-dummy interval segmentation provided by the TPEHGT dataset annotations, and (ii) fixed length blind segmentation into non overlapping 3-minute windows. In both cases, each segment is treated as one sample and inherits the record level label (term vs. preterm) of its originating pregnancy record.
II-C1 Contraction-dummy intervals
For the 2018 TPEHGT dataset, we use the contraction interval annotations provided with the dataset, together with the corresponding dummy (non contraction) intervals. Each annotated contraction interval and each dummy interval is treated as a separate sample, and the sample label is inherited from the mother/record label (term or preterm). Using the preterm subset yields $47$ contraction + $47$ dummy intervals ($94$ samples), while the term subset yields $53$ contraction + $53$ dummy intervals ($106$ samples). To address class imbalance for model training, term samples are randomly undersampled to match the $94$ preterm samples per iteration (details in Section II-F).
II-C2 Fixed 3-minute non overlapping intervals
To simulate a blind segmentation scenario without relying on contraction boundaries, each 30-minute record is partitioned into $10$ non overlapping windows of 3 minutes. We use 3-min windows because typical contractions last 1–2 min, and 3 min captures a full event plus baseline context. Each 3-minute window is treated as one sample and inherits the record label. For TPEHGT (excluding non pregnant records), this produces $26\times 10=260$ samples (130 term + 130 preterm). For TPEHG (300 records: 38 preterm, 262 term), this produces $300\times 10=3000$ windows in total (380 preterm + 2620 term). For balanced training iterations, all 380 preterm windows are used and 380 term windows are randomly selected per iteration.
II-D Feature Extraction
II-D1 Peak amplitude of the normalized power spectrum
The PA of the normalized power spectrum characterizes the dominant spectral component within a specified frequency band. This feature is designed to capture relative spectral concentration while remaining invariant to absolute signal amplitude.
@@ II-D1 Peak amplitude of the normalized power spectrum
For a discrete time signal $x[n]$, the power spectral density (PSD) $P(f_{i})$ is estimated using Welch’s method [18] over discrete frequency bins $f_{i}$. A frequency band $f_{\text{low}},f_{\text{high}}]$ is then selected, and the PSD values within this band are normalized by the total band power to ensure comparability across recordings of varying signal energy. The PA is defined as the maximum value of the band normalized PSD.
$A_{\text{peak}}=\max_{f_{i}\in[f_{\text{low}},f_{\text{high}}]}\frac{P(f_{i})}{\sum_{f_{j}\in[f_{\text{low}},f_{\text{high}}]}P(f_{j})}$ (8)
II-D2 Mel coefficients extraction
Mel Frequency Cepstral Coefficients (MFCCs) [19, 20] represent the short term power spectrum of a signal and are commonly used for feature extraction based on the Short-Time Fourier Transform (STFT). Initially developed for speech recognition, MFCCs have gained attention in biomedical signal processing due to their ability to capture perceptually relevant spectral features. MFCCs employ a Mel scale filter bank, which maps the frequency axis nonlinearly, providing higher resolution for lower frequencies. This emphasis is beneficial since both biomedical and acoustic signals are highly dependent on information present in the low frequency range [20].
@@ II-D2 Mel coefficients extraction
II-D3 Wavelet coefficients extraction
@@
In this work, the Daubechies 8 (db8) wavelet was selected because it effectively captures both the smooth low frequency uterine activity and the sharp transient changes, making it more suitable than shorter Daubechies wavelets. This property makes it particularly appropriate for analyzing the nonstationary nature of EHG signals [23]. Accordingly, the DWT was implemented using the db8 wavelet with a five level decomposition to analyze the EHG signals.
For each channel of the EHG signals, the five level wavelet decomposition produces five sets of detail coefficients (D1-D5) and one set of approximation coefficients (A5), yielding six sub bands in total. From each sub band, six statistical features were extracted from the wavelet coefficients to characterize the signal’s time-frequency properties. This results in 6 features $\times$ 6 sub bands, which is 36 features per channel. These features are computed from the coefficients within each sub band, where $c[i]$ represents the $i$-th wavelet coefficient and $N$ is the total number of coefficients in that sub band.
@@
II-E Model Selection
Given the characteristics of the dataset and the relatively limited size of available labeled EHG recordings, classical ML techniques were employed for classifying signal segments. These standard ML methods were chosen because, with limited data, more complex approaches such as DL can easily overfit. Coupled with the efficient denoising and feature selection strategies described in the previous sections, these computationally simpler methods are sufficient to capture relevant patterns. A set of supervised binary classification models was selected based on their proven performance in prior studies focused on PTB prediction and similar biomedical signal classification tasks [8, 7, 24]. The selected models include: